# PhishDebate: An LLM-Based Multi-Agent Framework for Phishing Website Detection

### v1

@@ V-D Case Study
PhishDebate begins each investigation by distributing the same artefacts —the raw URL [^1] The raw URL for this phishing website is: https://mail.mxcapital.com.br/wp-includes/wells/wells/page/index.htm., rendered HTML, and visible text—to its four specialised agents (URL Analyst, HTML Structure, Content Semantic, and Brand-Impersonation). Each agent reasons independently within its expertise, returns a claim (phishing or legitimate), a numerical confidence, and traceable evidence. The moderator then inspects the four perspectives and decides whether a clear majority exists. If not, it compels another debate round, asking every agent to reconcile its view with the counter-evidence provided by its peers. Only when consensus is reached (or the maximum number of rounds elapses) does the moderator submit a verdict to the judge, who finalises the assessment and records a confidence score.
Fig. 3 narrates a representative two-round session on a suspicious phishing webpage. In Round 1 the agents split 2–2: the URL Analyst and Brand-Impersonation agents flagged the directory string “/wp-includes/wells/wells/” as brand spoofing (Wells Fargo) and noted that legitimate sites never expose /wp-includes/; by contrast, the HTML Structure and Content Semantic agents found no typical phishing artefacts—no forms, iframes, or persuasive language—and therefore judged the page legitimate. Lacking consensus, the moderator declared the case UNCERTAIN and requested a second round.
@@
Figure 3: A detailed detection example with PhishDebate.
VI Conclusion
This paper presents PhishDebate, a modular multi-agent LLM-based debate framework for phishing website detection. By leveraging the collaborative reasoning of specialized agents and a structured debate process moderator and judged by dedicated agent, PhishDebate has demonstrated good performance on phishing website detection and addresses key limitations of traditional and single-agent LLM approaches.
Future research will explore the resilience of PhishDebate against adversarial phishing samples, including obfuscated, cloaked, or adversarially crafted webpages designed to bypass LLM-based detection. This will further validate the framework’s adaptability in real-world, adversarial threat environments.
@@ Appendix A Prompt Templates Used in This Study
- Confidence: [A score between 0 and 1]
- Evidence: [Relevant structural elements or tag patterns you found]
Figure 4: Prompt template for HTML Structure Agent.
- Claim: [Whether the page language seems phishing-related]
- Confidence: [A score between 0 and 1]
Figure 5: Prompt template for Content Semantic Agent.
Visible Text:
[TRUNCATED_VISIBLE_TEXT]
Provide response in following format:
- Claim: [Does the content attempt to impersonate a known brand?]
- Confidence: [A score between 0 and 1]
- Evidence: [Name(s) of impersonated brands and supporting context]
Figure 6: Prompt template for Brand Impersonation Agent.

### v2

@@ V-D Case Study
PhishDebate begins each investigation by distributing the same artefacts including the raw URL [^1] The raw URL for this phishing website is: https://mail.mxcapital.com.br/wp-includes/wells/wells/page/index.htm., rendered HTML, and visible text to its four specialised agents (URL Analyst, HTML Structure, Content Semantic, and Brand-Impersonation). Each agent reasons independently within its expertise, returns a claim (phishing or legitimate), a numerical confidence, and traceable evidence. The moderator then inspects the four perspectives and decides whether a clear majority exists. If not, it compels another debate round, asking every agent to reconcile its view with the counter-evidence provided by its peers. Only when consensus is reached (or the maximum number of rounds elapses) does the moderator submit a verdict to the judge, who finalises the assessment and records a confidence score.
Fig. 6 narrates a representative two-round session on a suspicious phishing webpage. In Round 1 the agents split 2–2: the URL Analyst and Brand-Impersonation agents flagged the directory string “/wp-includes/wells/wells/” as brand spoofing (Wells Fargo) and noted that legitimate sites never expose /wp-includes/; by contrast, the HTML Structure and Content Semantic agents found no typical phishing artefactssuch as forms, iframes, or persuasive language, and therefore judged the page legitimate. Lacking consensus, the moderator declared the case UNCERTAIN and requested a second round.
Figure 6: A detailed detection example with PhishDebate.
@@
VI Limitations
Despite its effectiveness, several limitations remain. First, the performance of debate agents ultimately reflects the biases and knowledge gaps of the underlying LLMs, which are bounded by their pre-training data. Second, input length restrictions of commercial LLM APIs constrain the amount of webpage content that can be fully processed, leading to truncation and potential information loss. Third, although our experiments employed commercial APIs, the core contribution of PhishDebate lies in demonstrating how debate-based reasoning mitigates hallucination and enhances interpretability. In practice, the same paradigm can be instantiated with locally deployed language models, enabling large-scale deployment under high-traffic conditions while simultaneously reducing cost and latency.
VII Conclusion and Future Directions
This paper introduced PhishDebate, a modular multi-agent LLM-based debate framework for phishing website detection. By leveraging collaborative reasoning among specialized agents under the coordination of a Moderator and a final Judge, PhishDebate achieved strong detection performance and addressed key limitations of traditional and single-agent LLM approaches, including limited interpretability and vulnerability to hallucination.
Future research will examine the resilience of PhishDebate against adversarial phishing tactics, such as obfuscation, cloaking, and adversarially crafted webpages, to further validate its adaptability in dynamic and hostile threat environments. Ultimately, we envision PhishDebate as an analyst-augmentation system that not only strengthens early, left-of-exploit detection but also scales to high-velocity, large-scale security data environments.